# Polynomial-order oscillations in geometric discrepancy

### v1

@@ Abstract
Let $C\subset\mathbb{R}^{2}$ be a convex body, and let $\mathcal{P}\subset[0,1)^{2}$ be a set of $N$ points. The discrepancy of $\mathcal{P}$ with respect to $C$ is defined as
$\mathcal{D}(\mathcal{P},\,C)=\sum_{\mathbf{p}\in\mathcal{P}}\sum_{\mathbf{n}\in\mathbb{Z}^{2}}\mathds{1}_{C}(\mathbf{p}+\mathbf{n})-N|C|.$
A standard approach for estimating how a point distribution deviates from uniformity is to average the latter quantity over a family of sets; in particular, when considering quadratic averages over translations and dilations, one obtains
$\mathcal{D}_{2}(\mathcal{P},\,C)=\int_{0}^{1}\int_{[0,1)^{2}}\left|\mathcal{D}(\mathcal{P},\,\boldsymbol{\tau}+\delta C)\right|^{2}\,\mathrm{d}\boldsymbol{\tau}\,\mathrm{d}\delta.$
This paper concerns the behaviour of optimal homothetic quadratic discrepancy
@@
Beck and Chen [MR1489133] showed that the optimal h.q.d. of convex polygons has an order of growth of $\log N$. More recently, Brandolini and Travaglini [MR4358540] showed that the optimal h.q.d. of planar convex bodies has an order of growth of $N^{1/2}$ if their boundary is $\mathcal{C}^{2}$, and of $N^{2/5}$ if their boundary is only piecewise-$\mathcal{C}^{2}$ and not polygonal. We show that, in general, no order of growth is required. First, by an implicit geometric construction, we show that one can obtain prescribed oscillations from a logarithmic order of $\log N$ to polynomial orders of $N^{2/5}$ and $N^{1/2}$, and vice versa. Secondly, by using Fourier-analytic methods, we show that prescribed polynomial-order oscillations in the range $N^{\alpha}$ with $\alpha\in(2/5,1/2)$ are achievable. These results provide the first examples of sets for which an optimal averaged discrepancy does not admit a single order of growth, and to the best of our knowledge, a similar phenomenon has not been reported before in the literature.
@@ 1 Introduction
The theory of irregularities of distribution, often referred to as discrepancy theory, quantifies how point distributions in a space deviate from uniformity (see [MR2683232] and [MR4391422] for a general overview and applications). In geometric discrepancy theory, historically initiated by Roth’s seminal work [MR66435], one measures such a difference using sets as test functions and compares the number of points that fall within them with the expected value. We introduce the following notions of discrepancy.
@@
Let $C\subset\mathbb{R}^{2}$ be a convex body, and let $\mathcal{P}_{N}\subset\mathbb{T}^{2}$ be a set of $N$ points. We define the discrepancy of $\mathcal{P}_{N}$ with respect to $C$ as
@@
Let $\boldsymbol{\tau}\in[0,1)^{2}$ be a translation parameter and $\delta\geq 0$ be a dilation parameter. We define the homothetic quadratic discrepancy (in short, h.q.d.) of $\mathcal{P}_{N}$ with respect to $C$ as
@@
It is a classic question in the theory of irregularities of distribution to determine the optimal behaviour of the h.q.d. as $N$ (the number of samples per unit square) goes to infinity. In other words, we are concerned with the quantity
@@
First, we state a result of Beck and Chen [MR1489133] on the optimal h.q.d. with respect to polygons.
@@
Secondly, we present a recent result of Brandolini and Travaglini [MR4358540] that, notably, recovers a polynomial order of growth across an interval of exponents.
@@
If $C$ is not a polygon and its boundary is piecewise-$\mathcal{C}^{2}$ with at least one corner, then it holds
$\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C)\approx_{C}N^{2/5}.$
Moreover, for $\alpha\in(2/5,1/2)$, there exists a planar convex body $C_{\alpha}$ such that
@@
In the current paper, we aim to show that, in general, the optimal h.q.d. does not need to exhibit an order of growth. This is achieved by presenting two general methods for constructing special sets whose optimal h.q.d. may continuously switch between different orders of growth (logarithmic or polynomial).
Our first method employs a simple geometric argument. This allows prescribed oscillations from a logarithmic order of $\log N$ to polynomial orders of $N^{2/5}$ and $N^{1/2}$, and vice versa. We state our result as follows.
@@
Let $\left\{\alpha_{i}\right\}_{i\in\mathbb{N}}$ be in $\{0,2/5,1/2\}$, let $\left\{\varepsilon_{i}\right\}_{i\in\mathbb{N}}$ be in $(0,\varepsilon]$ for some small positive $\varepsilon$, and let $\left\{q_{i}\right\}_{i\in\mathbb{N}}$ be positive integers. Then there exists a planar convex body $C$ and an increasing sequence $\left\{N_{i}\right\}_{i\in\mathbb{N}}\subset\mathbb{N}$ such that, for every $i\in\mathbb{N}$ such that $\alpha_{i}=0$, it holds
@@
while, for every $i\in\mathbb{N}$ such that $\alpha_{i}\in\{2/5,1/2\}$, it holds
This first method obtains the planar convex body $C$ as the limit of a nested sequence of planar convex bodies; thus, $C$ is constructed implicitly. The simplicity of our first method comes at the cost of limited knowledge on the geometry of the boundary of $C$.
@@
The second method involves constructing the boundary of $C$ “by hand”; as we will see, it suffices to design the boundary only in a small neighbourhood of a point, while leaving the remaining part as $\mathcal{C}^{2}$.
These two results settle the first examples of sets for which the optimal h.q.d. does not exhibit an order of growth. Interestingly, if one further considers rotations in the averaging process (therefore, averaging over the whole group of affine transformations), it is a classic result of Beck [MR0906524] that one always obtains an order of growth of $N^{1/2}$. The reader may consult a recent work by Gennaioli and the author [beretti2024fouriertransformbvfunctions] for further details on the interplay between discrepancy and geometric measure theory.
@@ 2 The first method
Recalling the hypotheses of Theorem 1.4, we have that $\left\{\alpha_{i}\right\}_{i\in\mathbb{N}}$ is a sequence of exponents with values in $\{0,2/5,1/2\}$, $\left\{\varepsilon_{i}\right\}_{i\in\mathbb{N}}$ is a (possibly decreasing) sequence in $(0,\varepsilon]$ for some small positive $\varepsilon$, and $\left\{q_{i}\right\}_{i\in\mathbb{N}}$ is a (possibly increasing) sequence of positive integers. We remark that $\alpha_{i}$ prescribes the target order of growth, while $\varepsilon_{i}$ and $q_{i}$ quantify, respectively, how closely and for how long we approximate it.
@@
Let $\mathcal{P}\subset\mathbb{T}^{2}$ be a set of $N$ points, and let $A$ and $B$ be two planar sets such that $\rm{diam}(A)\leq 1$, $A\supset B$, and $|A\setminus B|\leq\eta\leq 1$. Then, it holds
@@
By the Cauchy–Schwarz inequality, it holds
$\left|\mathcal{D}_{2}(\mathcal{P},A)-\mathcal{D}_{2}(\mathcal{P},B)\right|^{2}\leq 2N\int_{0}^{1}\int_{\mathbb{T}^{2}}\left|\mathcal{D}(\mathcal{P},\boldsymbol{\tau}+\delta A)-\mathcal{D}(\mathcal{P},\boldsymbol{\tau}+\delta B)\right|^{2}\,\mathrm{d}\boldsymbol{\tau}\,\mathrm{d}\delta.$ (2.1)
@@
has measure $\delta^{2}|A\setminus B|=\delta^{2}\eta$. Therefore, it holds
$\int_{\mathbb{T}^{2}}\left|\mathds{1}_{\boldsymbol{\tau}+\delta(A\setminus B)}(\mathbf{p})\right|^{2}\,\mathrm{d}\boldsymbol{\tau}=\delta^{2}\eta.$
Now, by applying the Cauchy–Schwarz inequality again, we obtain
@@
The latter result is all we need to proceed with the proof of Theorem 1.4. We start by considering a sequence of convex planar bodies $\left\{C_{i}\right\}_{i\in\mathbb{N}}$ to be chosen later, but such that each diameter is less than $1$ and
@@
$H_{i}(N)=\begin{cases}\log N&\text{if}\quad\alpha_{i}=0\\ N^{2/5}&\text{if}\quad\alpha_{i}=2/5\\ N^{1/2}&\text{if}\quad\alpha_{i}=1/2\\ \end{cases}.$
@@
We now construct the sequence $\left\{\eta_{i}\right\}_{i\in\mathbb{N}}$ recursively. Suppose $\left\{\eta_{j}\right\}_{j=1}^{i-1}$ and $\left\{C_{j}\right\}_{j=1}^{i-1}$ have already been chosen. Then, let $N_{i}$ be such that
@@
In particular, notice that such an $N_{i}$ exists, otherwise (2.3) would not hold. Then, choose $\eta_{i}$ in such a way that $\eta_{i}\leq\eta_{i-1}/2$ and
$\eta_{i}\leq(N_{i}+q_{i})^{-4}.$
By choosing the whole sequence $\left\{\eta_{i}\right\}_{i\in\mathbb{N}}$ in such a fashion, it follows that the nested sequence of the planar convex bodies $\left\{C_{i}\right\}_{i\in\mathbb{N}}$ converges (in measure) to a planar convex body $C$ such that, for every $i\in\mathbb{N}$, it holds
$|C\setminus C_{i}|\leq\sum_{j=i+1}^{\infty}\eta_{j}\leq\eta_{i}.$
Therefore, by applying Lemma 2.1, we obtain that, for every $i\in\mathbb{N}$, for every $N\in[N_{i},N_{i}+q_{i}]$, and for every set $\mathcal{P}\subset[0,1)^{2}$ of $N$ points, it holds
$|\mathcal{D}_{2}(\mathcal{P},\,C)-\mathcal{D}_{2}(\mathcal{P},\,C_{i})|\leq 2N^{2}\eta_{i}^{1/2}\leq 2.$
Finally, by (2.4) and a bit of rearrangement, the estimates in Theorem 1.4 follow.

### v2

@@ Abstract
Let $C\subset\mathbb{R}^{2}$ be a convex body, and for a positive integer $N$, let $\mathcal{P}$ be a configuration of $N$ points in $0,1)^{2}$. The discrepancy of $\mathcal{P}$ with respect to $C$ is defined by
$\mathcal{D}(\mathcal{P},\,C)=\sum_{\mathbf{p}\in\mathcal{P}}\sum_{\mathbf{n}\in\mathbb{Z}^{2}}\mathds{1}_{C}(\mathbf{p}+\mathbf{n})-N|C|,$
and one may estimate how $\mathcal{P}$ deviates from uniformity by averaging the latter quantity over a family of sets. When considering quadratic averages over translated and dilated copies of $C$, one gets the homothetic quadratic discrepancy
$\mathcal{D}_{2}(\mathcal{P},\,C)=\int_{0}^{1}\int_{[0,1)^{2}}\left|\mathcal{D}(\mathcal{P},\,\boldsymbol{\tau}+\delta C)\right|^{2}\,{\rm d}\boldsymbol{\tau}\,{\rm d}\delta.$
We investigate the behaviour of the optimal homothetic quadratic discrepancy, that is
@@
Beck [BEC88] and Beck and Chen [BC97] showed that the optimal h.q.d. of convex polygons has an order of growth of $\log N$, and more recently, Brandolini and Travaglini [BT22] proved that the optimal h.q.d. of planar convex bodies with a $\mathcal{C}^{2}$ boundary has an order of growth of $N^{1/2}$. We show that, in general, a single order of growth for the optimal h.q.d. need not exist. First, by an implicit geometric construction of $C$, we obtain prescribed oscillations between $\log N$ and $N^{1/2}$. Second, by a subtler design of $\partial C$ and via Fourier-analytic methods, we obtain prescribed polynomial-order oscillations in the range $N^{\alpha}$ with $\alpha\in(2/5,1/2)$. Moreover, we show that the set of planar convex bodies whose optimal h.q.d. does not admit a single order of growth is residual in the (Hausdorff) metric space of planar convex bodies.
@@ 1 Introduction
The theory of irregularities of distribution, often referred to as discrepancy theory, quantifies how point distributions in a space deviate from uniformity (see [MAT10] and [BDP20] for a general overview and applications). In geometric discrepancy theory, historically initiated by Roth’s seminal work [ROT54], one measures such a difference using sets as test functions and compares the number of points that fall within them with the expected value. We introduce the following notions of discrepancy.
@@
Let $C\subset\mathbb{R}^{2}$ be a convex body, and let $\mathcal{P}_{N}$ be a configuration of $N$ points (that is, a collection of $N$ not necessarily distinct points counted with multiplicity) in $\mathbb{T}^{2}$. We define the discrepancy of $\mathcal{P}_{N}$ with respect to $C$ as
@@
Let $\boldsymbol{\tau}\in[0,1)^{2}$ be a translation parameter and let $\delta\geq 0$ be a dilation parameter. We define the homothetic quadratic discrepancy (in short, h.q.d.) of $\mathcal{P}_{N}$ with respect to $C$ as
@@
It is a classical question in the theory of irregularities of distribution to determine the optimal behaviour of the h.q.d. as $N$ (the number of samples per unit square) goes to infinity. In other words, we are concerned with the quantity
@@
First, we state a result on the optimal h.q.d. with respect to polygons. Throughout the paper, we assume that $N$ is an integer greater than $1$.
@@
In particular, the lower bound is due to a classical work of Beck [BEC88], while the upper bound follows from a subsequent work of Beck and Chen [BC97].
In contrast with the latter result, one has the following result on the optimal h.q.d. for convex bodies with sufficiently regular boundaries.
@@
The upper bound is well-known and can be easily shown via a probabilistic argument employing a randomly shifted lattice (for example, see [BCC+19]). On the other hand, the lower bound is due to a recent work of Brandolini and Travaglini [BT22] and requires subtle Fourier-analytic techniques; these provide a starting point for our work. In the same paper, the following result is shown, notably recovering a polynomial order of growth across an interval of exponents.
Theorem 1.4 (Brandolini-Travaglini).
Let $\alpha\in[2/5,1/2)$. Then, there exists a planar convex body $C_{\alpha}$ such that
@@
In the current paper, we aim to show that, in general, the optimal h.q.d. does not need to exhibit a single order of growth, and prescribed stationary orders of growth can be achieved. For this purpose, we present two general methods for constructing special sets whose optimal h.q.d. continuously switch between different orders of growth (logarithmic or polynomial).
Our first method employs a simple geometric argument. This allows prescribed oscillations from a logarithmic order of $\log N$ to a polynomial order of $N^{1/2}$, and vice versa. We state our result as follows.
@@
Let $\left\{\alpha_{i}\right\}_{i\in\mathbb{N}}$ be in $\{0,1/2\}$, let $\left\{\varepsilon_{i}\right\}_{i\in\mathbb{N}}$ be in $(0,\varepsilon]$ for some small positive $\varepsilon$, and let $\left\{q_{i}\right\}_{i\in\mathbb{N}}$ be positive integers. Then there exists a planar convex body $C$ and an increasing sequence $\left\{N_{i}\right\}_{i\in\mathbb{N}}\subset\mathbb{N}$ such that, for every $i\in\mathbb{N}$ such that $\alpha_{i}=0$, it holds
@@
while, for every $i\in\mathbb{N}$ such that $\alpha_{i}=1/2$, it holds
$N^{1/2-\varepsilon_{i}}\leq\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C)\leq N^{1/2+\varepsilon_{i}}\quad\text{for every}\quad N\in[N_{i},N_{i}+q_{i}].$
This first method obtains the planar convex body $C$ as the limit of an increasing sequence of planar convex bodies; thus, $C$ is constructed implicitly. The simplicity of our first method comes at the cost of limited knowledge on the geometry of the boundary of $C$. In particular, we rely on Lemma 2.1, showing that two bodies with a small symmetric difference have close discrepancies. Further, by pairing the just-mentioned lemma with a classical topological argument, we show that the set of planar convex bodies whose optimal h.q.d. has a single order of growth is meagre in the space of planar convex bodies endowed with the Hausdorff metric.
@@
The second method involves constructing the boundary of $C$ “by hand”; as we will see, it suffices to design the boundary only in a small neighbourhood of a point, while leaving the remaining part as $\mathcal{C}^{2}$. Moreover, we note that we can alternate exponents within the same range as in Theorem 1.4.
To the best of our knowledge, these two results settle the first examples of sets for which the optimal h.q.d. has prescribed stationary orders of growth. Interestingly, if one further considers full rotations in the averaging process (that is, averaging over the whole similarity group), it is a classical result, obtained independently by Beck [BEC87] and Montgomery [MON94, Ch. 6], that one always obtains an order of growth of $N^{1/2}$. The reader may consult a recent work by Gennaioli and the author [BG24] for further details on the interplay between discrepancy and geometric measure theory.
@@ 2 The first method
Recalling the hypotheses of Theorem 1.5, we have that $\left\{\alpha_{i}\right\}_{i\in\mathbb{N}}$ is a sequence of exponents with values in $\{0,1/2\}$, $\left\{\varepsilon_{i}\right\}_{i\in\mathbb{N}}$ is a (possibly decreasing) sequence in $(0,\varepsilon]$ for some small positive $\varepsilon$, and $\left\{q_{i}\right\}_{i\in\mathbb{N}}$ is a (possibly increasing) sequence of positive integers. We remark that $\alpha_{i}$ prescribes the target order of growth, while $\varepsilon_{i}$ and $q_{i}$ quantify, respectively, how closely and for how long we approximate it.
@@
Let $\mathcal{P}$ be a configuration of $N$ points in $\mathbb{T}^{2}$, and let $A$ and $B$ be two planar sets such that $\rm{diam}(A)\leq 1$, $A\supset B$, and $|A\setminus B|\leq\eta\leq 1$. Then, it holds
@@
By the reverse triangle inequality, it holds
$\left|\mathcal{D}_{2}(\mathcal{P},A)-\mathcal{D}_{2}(\mathcal{P},B)\right|^{2}\leq 4N^{2}\int_{0}^{1}\int_{\mathbb{T}^{2}}\left|\mathcal{D}(\mathcal{P},\boldsymbol{\tau}+\delta A)-\mathcal{D}(\mathcal{P},\boldsymbol{\tau}+\delta B)\right|^{2}\,\mathrm{d}\boldsymbol{\tau}\,\mathrm{d}\delta.$ (2.1)
@@
has measure $\delta^{2}|A\setminus B|\leq\delta^{2}\eta$. Therefore, it holds
$\int_{\mathbb{T}^{2}}\left|\mathds{1}_{\boldsymbol{\tau}+\delta(A\setminus B)}(\mathbf{p})\right|^{2}\,\mathrm{d}\boldsymbol{\tau}\leq\delta^{2}\eta.$
Now, by applying the Cauchy–Schwarz inequality, we obtain
@@
The latter result is all we need to proceed with the proof of Theorem 1.5. We start by considering a sequence of convex planar bodies $\left\{C_{i}\right\}_{i\in\mathbb{N}}$ to be chosen later, but such that each of their diameters is less than $1$ and
@@
$H_{i}(N)=\begin{cases}\log N&\text{if}\quad\alpha_{i}=0\\ N^{1/2}&\text{if}\quad\alpha_{i}=1/2\end{cases}.$
@@
We now construct the sequence $\left\{\eta_{i}\right\}_{i\in\mathbb{N}}$ recursively. Suppose $\left\{\eta_{j}\right\}_{j=1}^{i}$ and $\left\{C_{j}\right\}_{j=1}^{i}$ have already been chosen. Then, let $N_{i}$ be such that
@@
In particular, notice that such an $N_{i}$ exists, or else it would contradict (2.3). Then, choose $\eta_{i+1}$ in such a way that $\eta_{i+1}\leq\eta_{i}/2$ and
$\eta_{i+1}<(N_{i}+q_{i})^{-4}/16.$
By choosing the whole sequence $\left\{\eta_{i}\right\}_{i\in\mathbb{N}}$ in such a fashion, it follows that the nested sequence of the planar convex bodies $\left\{C_{i}\right\}_{i\in\mathbb{N}}$ converges in the Hausdorff metric to a planar convex body $C$ such that, for every $i\in\mathbb{N}$, it holds
$|C\setminus C_{i}|\leq\sum_{j=i+1}^{\infty}\eta_{j}\leq 2\eta_{i+1}.$
Therefore, by applying Lemma 2.1, we obtain that, for every $i\in\mathbb{N}$, for every $N\in[N_{i},N_{i}+q_{i}]$, and for every configuration $\mathcal{P}$ of $N$ points in $\mathbb{T}^{2}$, it holds
$|\mathcal{D}_{2}(\mathcal{P},\,C)-\mathcal{D}_{2}(\mathcal{P},\,C_{i})|\leq 4N^{2}\eta_{i+1}^{1/2}<1.$
Finally, by pairing the latter inequality with (2.4), the claim of Theorem 1.5 follows.
2.1 The discrepancy of most convex bodies
What follows is inspired by [BRT98, Sec. 4], where the authors applied an analogous argument to study the average decay of the Fourier transform of convex sets. We let $\mathfrak{C}$ be the space of convex bodies in $\mathbb{R}^{2}$ endowed with the Hausdorff metric $d_{H}$ defined by
$d_{H}(A,B)=\max\Big\{\sup_{x\in A}\inf_{y\in B}|x-y|,\,\sup_{y\in B}\inf_{x\in A}|x-y|\Big\}\quad\text{for }A,B\in\mathfrak{C}.$
By the Blaschke selection theorem, the metric space $(\mathfrak{C},d_{H})$ is locally compact, and by the Baire category theorem, it follows that $(\mathfrak{C},d_{H})$ is a Baire space and, in particular, it is nonmeagre.
In [SW65] it is shown that the Hausdorff metric and the symmetric-difference metric $d_{\Delta}(A,B)=|A\Delta B|$ induce the same topology on $\mathfrak{C}$. It is not difficult to adapt the proof of Lemma 2.1 and show that, for a fixed $N\in\mathbb{N}$, the functional
$\mathcal{D}_{2}(\mathcal{P},\,\cdot)\colon(\mathfrak{C},d_{\Delta})\to(0,+\infty)$
is (locally) uniformly continuous in its argument with respect to all configurations $\mathcal{P}$ of $N$ points. Indeed, it is enough to pair the equality $A\Delta B=(A\setminus B)\sqcup(B\setminus A)$ with the fact that $\mathcal{D}(\mathcal{P},\,\cdot)$ is additive on disjoint sets, and note that the condition on the diameter in the hypothesis of Lemma 2.1 is (locally) not restrictive. Hence, we infer that, for a fixed $N\in\mathbb{N}$, the functional
$\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,\cdot)\colon(\mathfrak{C},d_{\Delta})\to(0,+\infty)$
is continuous. It is time to state a technical result of Gruber [GRU83].
Lemma 2.2 (Gruber).
Let $\{a_{j}\}_{j\in\mathbb{N}}\subset(0,+\infty)$, let $T$ be a nonmeagre topological space, and let $\{\phi_{j}\}_{j\in\mathbb{N}}$ be a sequence of continuous functionals with $\phi_{j}\,\colon\,T\to(0,+\infty)$. Define
$\mathcal{A}$ $=\big\{x\in T\;\colon\;\phi_{j}(x)=o(a_{j})\,\textnormal{ as }\,j\to+\infty\big\},$
$\mathcal{B}$ $=\big\{x\in T\;\colon\;a_{j}=o(\phi_{j}(x))\,\textnormal{ as }\,j\to+\infty\big\}.$
Then, the following holds:
  1. i)
    If $\mathcal{A}$ is dense in $T$, then, for every $x\in T$ except for a meagre subset, the inequality $\phi_{j}(x)<a_{j}$ holds for infinitely many values of $j$.
  2. ii)
    If $\mathcal{B}$ is dense in $T$, then, for every $x\in T$ except for a meagre subset, the inequality $\phi_{j}(x)>a_{j}$ holds for infinitely many values of $j$.
Finally, by pairing Theorem 1.2 and Theorem 1.3 with the latter lemma and the fact that $\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,\cdot)$ is continuous in $(\mathfrak{C},d_{\Delta})$, we obtain the following result.
Theorem 2.3.
Let $\varepsilon\in(0,1]$. Then, for every planar convex body $C\in\mathfrak{C}$ except for a meagre subset, there exist two infinite subsets $\mathcal{N}^{-}_{C},\mathcal{N}^{+}_{C}\subset\mathbb{N}$ such that
$\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C)\leq\log^{1+\varepsilon}N\quad\forall N\in\mathcal{N}^{-}_{C}\quad\textnormal{ and }\quad\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C)\geq N^{1/2-\varepsilon}\quad\forall N\in\mathcal{N}^{+}_{C}$.